OpenStudy (anonymous):

what is the equation of the line in standard form that passes throught hte point (3 4) and is parallel to the line y=-x+6 a) x+y=1 b) x+y=7 c) x-y=-1 d) x-y=-7

OpenStudy (anonymous):

please help !!!

OpenStudy (anonymous):

b

OpenStudy (anonymous):

The parallel line has the same slope as the given line so is of form y=-x+c. You know that the parallel line passes through (3,4) so substituting in gives 3=-4+c so c= 7 So you have y = -x+7 which is the same as b).